\begin{thebibliography}{22}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[mar()]{mardia_directional_1999}
Directional statistics.
\newblock URL \url{http://doi.wiley.com/10.1002/9780470316979}.

\bibitem[Banerjee et~al.(2005)Banerjee, Dhillon, Ghosh, and
  Sra]{JMLR:v6:banerjee05a}
Arindam Banerjee, Inderjit~S. Dhillon, Joydeep Ghosh, and Suvrit Sra.
\newblock Clustering on the unit hypersphere using von mises-fisher
  distributions.
\newblock \emph{Journal of Machine Learning Research}, 6\penalty0
  (46):\penalty0 1345--1382, 2005.
\newblock URL \url{http://jmlr.org/papers/v6/banerjee05a.html}.

\bibitem[Buzzega et~al.(2020)Buzzega, Boschini, Porrello, Abati, and
  CALDERARA]{NEURIPS2020_b704ea2c}
Pietro Buzzega, Matteo Boschini, Angelo Porrello, Davide Abati, and SIMONE
  CALDERARA.
\newblock Dark experience for general continual learning: a strong, simple
  baseline.
\newblock In H.~Larochelle, M.~Ranzato, R.~Hadsell, M.F. Balcan, and H.~Lin
  (eds.), \emph{Advances in Neural Information Processing Systems}, volume~33,
  pp.\  15920--15930. Curran Associates, Inc., 2020.
\newblock URL
  \url{https://proceedings.neurips.cc/paper/2020/file/b704ea2c39778f07c617f6b7ce480e9e-Paper.pdf}.

\bibitem[Cha et~al.(2021)Cha, Lee, and Shin]{Cha_2021_ICCV}
Hyuntak Cha, Jaeho Lee, and Jinwoo Shin.
\newblock Co2l: Contrastive continual learning.
\newblock In \emph{Proceedings of the IEEE/CVF International Conference on
  Computer Vision (ICCV)}, pp.\  9516--9525, 2021.

\bibitem[Chaudhry et~al.(2020)Chaudhry, Gordo, Lopez-Paz, Dokania, and
  Torr]{chaudhry2020using}
Arslan Chaudhry, Albert Gordo, David Lopez-Paz, Puneet~K. Dokania, and Philip
  Torr.
\newblock Using hindsight to anchor past knowledge in continual learning, 2020.
\newblock URL \url{https://openreview.net/forum?id=Hke12T4KPS}.

\bibitem[Chen et~al.(2020)Chen, Kornblith, Norouzi, and Hinton]{simclr}
Ting Chen, Simon Kornblith, Mohammad Norouzi, and Geoffrey Hinton.
\newblock A simple framework for contrastive learning of visual
  representations.
\newblock In Hal~Daumé III and Aarti Singh (eds.), \emph{Proceedings of the
  37th International Conference on Machine Learning}, volume 119 of
  \emph{Proceedings of Machine Learning Research}, pp.\  1597--1607. PMLR, Jul
  2020.
\newblock URL \url{https://proceedings.mlr.press/v119/chen20j.html}.

\bibitem[Davari et~al.(2022)Davari, Asadi, Mudur, Aljundi, and
  Belilovsky]{Davari_2022_CVPR}
MohammadReza Davari, Nader Asadi, Sudhir Mudur, Rahaf Aljundi, and Eugene
  Belilovsky.
\newblock Probing representation forgetting in supervised and unsupervised
  continual learning.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition (CVPR)}, pp.\  16712--16721, June 2022.

\bibitem[Gallardo et~al.(2021)Gallardo, Hayes, and
  Kanan]{Gallardo2021SelfSupervisedTE}
Jhair Gallardo, Tyler~L. Hayes, and Christopher Kanan.
\newblock Self-supervised training enhances online continual learning.
\newblock In \emph{British Machine Vision Conference}, 2021.

\bibitem[Gopal \& Yang(2014)Gopal and Yang]{pmlr-v32-gopal14}
Siddharth Gopal and Yiming Yang.
\newblock Von mises-fisher clustering models.
\newblock In Eric~P. Xing and Tony Jebara (eds.), \emph{Proceedings of the 31st
  International Conference on Machine Learning}, volume~32 of \emph{Proceedings
  of Machine Learning Research}, pp.\  154--162, Bejing, China, 22--24 Jun
  2014. PMLR.
\newblock URL \url{https://proceedings.mlr.press/v32/gopal14.html}.

\bibitem[Han \& Guo(2021)Han and Guo]{han2021contrastive}
Xuejun Han and Yuhong Guo.
\newblock Contrastive continual learning with feature propagation.
\newblock \emph{arXiv preprint arXiv:2112.01713}, 2021.

\bibitem[Hasnat et~al.(2017)Hasnat, Bohn{\'e}, Milgram, Gentric, and
  Chen]{Hasnat2017vonMM}
Md.~Abul Hasnat, Julien Bohn{\'e}, Jonathan Milgram, St{\'e}phane Gentric, and
  Liming Chen.
\newblock von mises-fisher mixture model-based deep learning: Application to
  face verification.
\newblock \emph{ArXiv}, abs/1706.04264, 2017.

\bibitem[He et~al.(2016)He, Zhang, Ren, and Sun]{7780459}
K.~He, X.~Zhang, S.~Ren, and J.~Sun.
\newblock Deep residual learning for image recognition.
\newblock In \emph{2016 IEEE Conference on Computer Vision and Pattern
  Recognition (CVPR)}, pp.\  770--778, Los Alamitos, CA, USA, jun 2016. IEEE
  Computer Society.
\newblock \doi{10.1109/CVPR.2016.90}.
\newblock URL \url{https://doi.ieeecomputersociety.org/10.1109/CVPR.2016.90}.

\bibitem[Hwang et~al.(2019)Hwang, Yu, Shi, Collins, Yang, Zhang, and
  Chen]{hwang2019segsort}
Jyh-Jing Hwang, Stella~X Yu, Jianbo Shi, Maxwell~D Collins, Tien-Ju Yang, Xiao
  Zhang, and Liang-Chieh Chen.
\newblock Segsort: Segmentation by discriminative sorting of segments.
\newblock In \emph{Proceedings of the IEEE/CVF International Conference on
  Computer Vision}, pp.\  7334--7344, 2019.

\bibitem[Khosla et~al.(2020)Khosla, Teterwak, Wang, Sarna, Tian, Isola,
  Maschinot, Liu, and Krishnan]{NEURIPS2020_d89a66c7}
Prannay Khosla, Piotr Teterwak, Chen Wang, Aaron Sarna, Yonglong Tian, Phillip
  Isola, Aaron Maschinot, Ce~Liu, and Dilip Krishnan.
\newblock Supervised contrastive learning.
\newblock In H.~Larochelle, M.~Ranzato, R.~Hadsell, M.F. Balcan, and H.~Lin
  (eds.), \emph{Advances in Neural Information Processing Systems}, volume~33,
  pp.\  18661--18673. Curran Associates, Inc., 2020.
\newblock URL
  \url{https://proceedings.neurips.cc/paper/2020/file/d89a66c7c80a29b1bdbab0f2a1a94af8-Paper.pdf}.

\bibitem[Kobayashi(2021)]{Kobayashi_2021_CVPR}
Takumi Kobayashi.
\newblock T-vmf similarity for regularizing intra-class feature distribution.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition (CVPR)}, pp.\  6616--6625, June 2021.

\bibitem[Krizhevsky et~al.(2009)Krizhevsky, Hinton,
  et~al.]{krizhevsky2009learning}
Alex Krizhevsky, Geoffrey Hinton, et~al.
\newblock Learning multiple layers of features from tiny images.
\newblock 2009.

\bibitem[Mai et~al.(2021{\natexlab{a}})Mai, Li, Kim, and Sanner]{Mai_2021_CVPR}
Zheda Mai, Ruiwen Li, Hyunwoo Kim, and Scott Sanner.
\newblock Supervised contrastive replay: Revisiting the nearest class mean
  classifier in online class-incremental continual learning.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision
  and Pattern Recognition (CVPR) Workshops}, pp.\  3589--3599, June
  2021{\natexlab{a}}.

\bibitem[Mai et~al.(2021{\natexlab{b}})Mai, Li, Kim, and
  Sanner]{Mai2021SupervisedCR}
Zheda Mai, Ruiwen Li, Hyunwoo~J. Kim, and Scott Sanner.
\newblock Supervised contrastive replay: Revisiting the nearest class mean
  classifier in online class-incremental continual learning.
\newblock \emph{2021 IEEE/CVF Conference on Computer Vision and Pattern
  Recognition Workshops (CVPRW)}, pp.\  3584--3594, 2021{\natexlab{b}}.

\bibitem[Van~der Maaten \& Hinton(2008)Van~der Maaten and
  Hinton]{van2008visualizing}
Laurens Van~der Maaten and Geoffrey Hinton.
\newblock Visualizing data using t-sne.
\newblock \emph{Journal of machine learning research}, 9\penalty0 (11), 2008.

\bibitem[Wang et~al.(2023)Wang, Wu, Chen, Liu, Fu, Huang, Zhou, Zhan, and
  Hua]{wang2023deep}
Pengfei Wang, Daqing Wu, Chong Chen, Kunpeng Liu, Yanjie Fu, Jianqiang Huang,
  Yuanchun Zhou, Jianfeng Zhan, and Xiansheng Hua.
\newblock Deep adaptive graph clustering via von mises-fisher distributions.
\newblock \emph{ACM Transactions on the Web}, 2023.

\bibitem[Zhang et~al.(2020)Zhang, Shen, and
  Deng]{DBLP:journals/corr/abs-2006-05882}
Song Zhang, Gehui Shen, and Zhi{-}Hong Deng.
\newblock Self-supervised learning aided class-incremental lifelong learning.
\newblock \emph{CoRR}, abs/2006.05882, 2020.
\newblock URL \url{https://arxiv.org/abs/2006.05882}.

\bibitem[Zhe et~al.(2019)Zhe, Chen, and Yan]{ZHE2019113}
Xuefei Zhe, Shifeng Chen, and Hong Yan.
\newblock Directional statistics-based deep metric learning for image
  classification and retrieval.
\newblock \emph{Pattern Recognition}, 93:\penalty0 113--123, 2019.
\newblock ISSN 0031-3203.
\newblock \doi{https://doi.org/10.1016/j.patcog.2019.04.005}.
\newblock URL
  \url{https://www.sciencedirect.com/science/article/pii/S0031320319301451}.

\end{thebibliography}

\documentclass{article} 
\usepackage{iclr2023_conference_tinypaper,times}

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}

\def\1{\bm{1}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\usepackage{hyperref}
\usepackage{url}
\usepackage{multirow}

\title{Towards Robust Feature Learning with t-vFM Similarity for Continual Learning}

\author{Bilan Gao \\
Department of Artificial Intelligent\\
Chung-Ang University\\
Seoul, Korea \\
\texttt{bilan2020@cau.ac.kr} \\
\And
YoungBin Kim \\
Department of Imaging Science and Arts \\
The Graduate School of Advanced Imaging Science,\\ Multimedia and Film,\\
Chung-Ang University \\
Seoul, Korea \\
\texttt{ybkim85@cau.ac.kr} \\
}

\iclrfinalcopy  
\begin{document}

\maketitle

\begin{abstract}
Continual learning has been developed using standard supervised contrastive loss from the perspective of feature learning. Due to the data imbalance during the training, there are still challenges in learning better representations. In this work, we suggest using a different similarity metric instead of cosine similarity in supervised contrastive loss in order to learn more robust representations. We validate the our method on one of the image classification datasets Seq-CIFAR-10 and the results outperform recent continual learning baselines.
\end{abstract}
\section{Introduction}
Continual learning is the research direction that mainly tackles the problem of catastrophic forgetting while a model learns new data consistently. To mitigate the catastrophic forgetting phenomenon, one strategy is to enhance the learned representation while training. An effective approach for achieving this is to apply supervised contrastive learning \citep{Cha_2021_ICCV,Mai_2021_CVPR,Mai2021SupervisedCR,han2021contrastive,Davari_2022_CVPR}. Neglecting the consideration of the imbalance phenomenon between current training samples and replayed samples can cause loss of  intra-class information while learning representations.This is because cosine similarity within the supervised contrastive loss, can extract features with large within-class variance, resulting in degraded model performance. To enhance the stability and reliability of feature extraction, we drew inspiration from \citep{Kobayashi_2021_CVPR}, we propose to use a similarity function based on the von Mises-Fisher distribution, further extended with a student t-distribution(t-vMF) to replace cosine similarity.  We adopt the method on asymmetric supervised contrastive loss\citep{Cha_2021_ICCV}, and conduct experiments under offline continual learning settings, and the results show effective improvement on prediction accuracy. 

\section{t-vMF Similarity Adopted in Continual Representation Learning}
An overview of asymmetric supervised contrastive loss ${L}_{ASC}$\citep{Cha_2021_ICCV}. A base encoder $f(\cdot)$ receives as input two augmented versions $\tilde{x}_{i}$ and $\tilde{x}_{p}$ from data sample $x$ with it's label $y$.  To using an inner product to measure distances, this representations are further mapped to a feature space with projection model $g(\cdot)$. The final extracted features can be written as $z_{i}=g(f(\tilde{x}_{i}))$ and $z_{p}=g(f(\tilde{x}_{p}))$. Differ from \citep{Mai_2021_CVPR}, ${L}_{ASC}$ selects anchor samples only from current training data rather than consider all the samples in the training stage. Let $S$ denotes all the indices of the current learning samples in a single batch and $P(i)$ denotes the indices of all positive samples that are distinguished from $i$. ${L}_{ASC}$ can be defined as follows:
\begin{equation}\label{eq_1}
    {L}_{ASC}={\sum_{i\in S}\frac{-1}{|P(i)|}{\sum_{p\in P(i)} {log}\frac{exp(sim({z}_{i}\cdot {z}_{p})/{\tau})}{\sum_{a\in A(i)}exp(sim({z}_{i}\cdot {z}_{a})/{\tau})})}},
\end{equation}
where $sim({z}_{i}\cdot {z}_{p})$, a cosine similarity function can be written as $= \|{z}_{i}\|\|{z}_{p}\|cos\theta$ with an angle $\theta$ between ${z}_{i}$ and ${z}_{p}$. And ${\tau}$ is a temperature hyperparameter. $P(i)$ denotes the indices of all positive samples that are distinguish from i; it is equivalent to $\{p\in A(i): \tilde{y}_{p}=\tilde{y}_{i}\}$. $|P(i)|$ is the number of elements in set $P(i)$.\\
We propose to use the t-vMF Similarity\citep{Kobayashi_2021_CVPR} to replace the cosine similarity in equation\ref{eq_1}. We first build a von Mises–Fisher distribution\citep{mardia_directional_1999,JMLR:v6:banerjee05a} based similarity $p(\Tilde{z}_{i};\Tilde{z}_{p},\kappa)={C}_{\kappa}exp(\kappa cos\theta)$ between features $\Tilde{z}_{i}$ and $\Tilde{z}_{p}$. It produces a probability density function on $p$ dimension with with parameter $\kappa$. The parameter $\kappa$ controls 
concentration of the distribution, and ${C}_{\kappa}$ is a normalization constant. Apply a t-distributed profile function $f_{t}(d;\kappa)=\frac{1}{1+\frac{1}{2}\kappa{d}^{2}}$ on $p(\Tilde{z}_{i};\Tilde{z}_{p},\kappa)$, the equation of t-vFM similarity function between two features $\Tilde{z}_{i}$ and $\Tilde{z}_{p}$ is formalized as follows(More detail in Appendix \ref{appendix_1}):
\begin{equation}\label{eq_2}
    \phi_{t}(cos\theta;\kappa)==\frac{1+cos{\theta}}{1+\kappa(1-cos{\theta})}-1
\end{equation}

Adopt equation\ref{eq_2} in the original ${L}_{ASC}$, the loss function of our proposed method is shown as follows:
\begin{equation}\label{eq_3}
    {L}_{Ours}={\sum_{i\in S}\frac{-1}{|P(i)|}{\sum_{p\in P(i)} {log}\frac{exp(\phi_{t}(cos\theta;\kappa)/{\tau})}{\sum_{a\in A(i)}exp(\phi_{t}(cos\theta;\kappa)/{\tau})})}}.
\end{equation}
The larger value $\kappa$ has, that smaller compact region between two representation is.
\\
\begin{table}[t]
\centering
\caption{Classification accuracy for Seq-CIFAR-10. All results are
averaged over ten independent trials. Best performance are marked in bold.}
\label{table1}
\begin{center}
\begin{tabular}{llll}
\multicolumn{1}{c}{\bf Buffer}  &\multicolumn{1}{c}{\bf Baselines} &\multicolumn{1}{c}{\bf Class Incremental} &\multicolumn{1}{c}{\bf Task Incremental}
\\ \hline \\
\multirow{4}{4em}{200} &$HAL$ \citep{chaudhry2020using}       &$32.36\pm 2.70$  &$82.51\pm 3.20$\\
             &$DER$ \citep{NEURIPS2020_b704ea2c}    &$61.93\pm 1.79$  &$91.40\pm 0.92$\\
             &$DER++$ \citep{NEURIPS2020_b704ea2c}  &$64.88\pm 1.17$  &$91.92\pm 0.60$\\
             &$Co^{2}L$ \citep{Cha_2021_ICCV}     &$65.57\pm 1.37$  &$93.43\pm 0.78$\\
             &\textbf{Ours}                     &$\textbf{67.98}\pm \textbf{1.29}$  &$\textbf{95.10}\pm \textbf{0.42}$\\
\\ \hline \\
\multirow{4}{4em}{500} &$HAL$\citep{chaudhry2020using}       &$41.79\pm 4.46$  &$84.54\pm 2.36$\\
             &$DER$ \citep{NEURIPS2020_b704ea2c}    &$70.51\pm 1.67$  &$93.40\pm 0.39$\\
             &$DER++$ \citep{NEURIPS2020_b704ea2c}  &$72.70\pm 1.17$  &$93.88\pm 0.60$\\
             &$Co^{2}L$ \citep{Cha_2021_ICCV}     &$\textbf{74.26}\pm \textbf{0.77}$  &$\textbf{95.90}\pm \textbf{0.26}$\\
             &\textbf{Ours}                     &$71.68\pm 0.41$  &$95.51\pm 0.19$\\
\\ \hline \\
\end{tabular}
\end{center}
\end{table}
\section{Experiment and Discussion}
Following prior work \citep{Cha_2021_ICCV}, we demonstrate the our method by training ResNet-18 \citep{7780459} as a base encoder on Seq-CIFAR-10 dataset \citep{krizhevsky2009learning}. We validate our method on class incremental learning and task incremental learning scenarios with replay buffers of size 200 and 500 \citep{NEURIPS2020_b704ea2c}.The value of $\kappa$ in Equation \ref{eq_3} is set to 16.(More detail in Appendix \ref{appendix_2}). 
As shown in Table \ref{table1}, our proposed method outperforms all the baselines with buffer 200, which indicates that the t-vMF similarity shows promise. In the case of buffer size 500, even though the proposed method does not outperform all baselines, it approaches the best baselines.
\section{Conclusion and Future Work}
In this work, we delve deeply into continual representation learning, using t-vMF similarity to enhance feature reliability and outperform most baselines in terms of prediction accuracy. Our future aim is to further investigate alternative representation learning approaches and various metrics to improve continual representation learning.

\subsubsection*{Acknowledgements}
This research was supported by Basic Science Research Program through the National Research Foundation of Korea(NRF) funded by the Ministry of Education(NRF-2022R1C1C1008534), and Institute for Information \& communications Technology Planning \& Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program, Chung-Ang University).

\subsubsection*{URM Statement}
Author Bilan Gao meets the URM criteria of ICLR 2023 Tiny Papers Track.

\bibliography{iclr2023_conference_tinypaper}
\bibliographystyle{iclr2023_conference_tinypaper}

\appendix
\section{Appendix}
\subsection{More detail explanation on T-vFM similarity}\label{appendix_1}
T-vFM similarity has the foundation on a similarity function that based von Mises–Fisher distribution\citep{mardia_directional_1999,JMLR:v6:banerjee05a}, a von Mises–Fisher distributed similarity which produces a probability density function on $p$ dimension with parameter $\kappa$ between two features $\Tilde{z}_{i}$ and $\Tilde{z}_{p}$ in $d$ dimension can be formulated as:
\begin{align}
  p(\Tilde{z}_{i};\Tilde{z}_{p},\kappa)& ={C}_{\kappa}exp(\kappa\Tilde{z}_{p}^\top\Tilde{z}_{i})\\
    & ={C}_{\kappa}exp(\kappa cos\theta).
\label{eq_5}   
\end{align}
The parameter $\kappa$ controls concentration of the distribution, and ${C}_{\kappa}$ is a normalization constant. To find out the part that mainly support the distribution $p(\Tilde{z}_{i};\Tilde{z}_{p},\kappa)$, define a profile function $f_{e}$ with $d$ and $\kappa$, then the profile function $f_{e}(d;\kappa)$ is defined as
\begin{equation}
    f_{e}(d;\kappa)=exp(-\frac{1}{2}\kappa{d}^{2}).
    \label{eq_6}
\end{equation}
Thus, Equation \ref{eq_5} can be reformulated by using a profile function(Equation \ref{eq_6}) as follows
\begin{align}
    p(\Tilde{z}_{i};\Tilde{z}_{p},\kappa)& ={C}_{\kappa}exp(\kappa-\frac{1}{2}\kappa\|\Tilde{z}_{i}-\Tilde{z}_{p}\|^{2})\\
    & ={C}^{'}_{\kappa}f_{e}(\|\Tilde{z}_{i}-\Tilde{z}_{p}\|;\kappa).
\label{eq_8}
\end{align}
Observing from Equation \ref{eq_8}, the vMF similarity is primarily distinguished by $f_{e}(\|\Tilde{z}_{i}-\Tilde{z}_{p}\|;\kappa)$, which means the vMF similarity can formally be re-defined as an alternative for $cos\theta$ as
\begin{align}
\phi_{e}(cos\theta;\kappa)& =2\frac{f_{e}(|\Tilde{z}_{i}-\Tilde{z}_{p}\|;\kappa)-f_{e}(2;\kappa)}{f_{e}(0;\kappa)-f_{e}(2;\kappa)}-1\\
& =2\frac{exp({\kappa}cos\theta)-exp(-\kappa)}{exp(\kappa)-exp(-\kappa)}-1 \in [1,-1].
\label{eq_10}
\end{align}
Equation \ref{eq_10} demonstrates von Mises–Fisher distributed similarity between two features $\Tilde{z}_{i}$ and $\Tilde{z}_{p}$ with an important parameter $\kappa$. According to its function that introduced above, changing different value of $\kappa$ can dynamically balance the trade-off between learning discriminative features for each task and minimizing interference between new task samples and learned task samples.\\ 
A previous work \cite{van2008visualizing} suggested using a heavy tail distribution to better capture distinguishing features. Modify the profile function(Equation \ref{eq_6}) with a heavy-tailed student T distribution, the new profile function can be updated as follows:
\begin{equation}
    f_{t}(d;\kappa)=\frac{1}{1+\frac{1}{2}\kappa{d}^{2}}.
    \label{eq_11}
\end{equation}
Due to the property of t-distribution, it makes the similarity function less sensitive to small changes in the learned representations. We reformulate vMF similarity with Equation \ref{eq_10}, the formal definition of t-vFM similarity between features $\Tilde{z}_{i}$ and $\Tilde{z}_{p}$ is given as follows:
\begin{align}
\phi_{t}(cos\theta;\kappa)& =2\frac{f_{t}(|\Tilde{z}_{i}-\Tilde{z}_{p}\|;\kappa)-f_{t}(2;\kappa)}{f_{t}(0;\kappa)-f_{t}(2;\kappa)}-1,\\
& =2\frac{\frac{1}{1+\kappa(1-cos{\theta})-\frac{1}{1+2\kappa}}}{1-\frac{1}{1+2\kappa}}-1,\\
&=\frac{1+cos{\theta}}{1+\kappa(1-cos{\theta})}-1.
\label{eq_14}
\end{align}
As shown in Equation \ref{eq_14}, t-vFM similarity can be simply calculated with $cos\theta$ and $\kappa$. Since $cos\theta \in [-1,+1]$, $f_{e}(\|\Tilde{z}_{i}-\Tilde{z}_{p}\|;\kappa)$ is resized to be adaptable with $cos\theta$. Applying Equation \ref{eq_14} to asymmetric contrastive loss, our proposed loss can be modified from Equation \ref{eq_1} as follows:
\begin{equation}
    {L}_{Ours}={\sum_{i\in S}\frac{-1}{|P(i)|}{\sum_{p\in P(i)} {log}\frac{exp(\phi_{t}(cos\theta;\kappa)/{\tau})}{\sum_{a\in A(i)}exp(\phi_{t}(cos\theta;\kappa)/{\tau})})}}.
    \label{eq14}
\end{equation}

\subsection{Experimental Details}\label{appendix_2}
\paragraph{Dataset and architecture}
In the training stage, we split the CIFAR-10 dataset into five distinct sets of samples,with each set consisting of two different classes. For architecture, we use a non-pretrained ResNet-18 as a base encoder for extracting features followed by a 2-layer projection MLP to map these representations to a latent space of 128 dimensions \citep{NEURIPS2020_d89a66c7}.
\paragraph{Hyperparameters} We train the base encoder with a 512 batch size, and 0.5 as learning rate. The temperature hyperparameter($\tau$) for our loss is set to 0.5.
\paragraph{Evaluation}We follow class balance strategy\cite{Cha_2021_ICCV} and use an additional linear classifier trained only on the final task samples and the buffer data with learned representations. All the results are reported with 10 trials.
\paragraph{Additional experiment on different values of $\kappa$}
As shown in Table \ref{table2}, we report results with different $\kappa$ values, and our algorithm surpasses all the baselines of recent works in the case of buffer 200 setting. In the comparison of 3 different values of $\kappa$, class incremental learning reaches the best performance when $\kappa=32$, and task incremental learning achieves the best performance when $\kappa=4$. Such results indicate that our method successfully learns and reduces the effect of data imbalance. Moreover, there are no big difference among all the results with different values of $\kappa$, which indicates a narrow region does not effect the performance in this case.
\begin{table}[t]
\centering
\caption{Classification accuracy for Seq-CIFAR-10. All results are
averaged over ten independent trials. Best performance are marked in bold. Our proposed method choose three values of $\kappa$: 4,16,32}
\label{table2}
\begin{center}
\begin{tabular}{llll}
\multicolumn{1}{c}{\bf Buffer}  &\multicolumn{1}{c}{\bf Baselines} &\multicolumn{1}{c}{\bf Class Incremental} &\multicolumn{1}{c}{\bf Task Incremental}
\\ \hline \\
\multirow{7}{4em}{200} &$HAL$ \citep{chaudhry2020using}       &$32.36\pm 2.70$  &$82.51\pm 3.20$\\
             &$DER$ \citep{NEURIPS2020_b704ea2c}    &$61.93\pm 1.79$  &$91.40\pm 0.92$\\
             &$DER++$ \citep{NEURIPS2020_b704ea2c}  &$64.88\pm 1.17$  &$91.92\pm 0.60$\\
             &$Co^{2}L$ \citep{Cha_2021_ICCV}     &$65.57\pm 1.37$  &$93.43\pm 0.78$\\
             &$Ours(\kappa=4)$ &$67.94\pm 1.04$  
             &$\textbf{95.20}\pm \textbf{0.22}$\\
             
             &$Ours$ $(\kappa =16)$                  &$67.98\pm 1.29$  &$95.10\pm 0.42$\\
             &$Ours$ $(\kappa =32)$                  &$\textbf{67.99}\pm \textbf{0.71}$  &$95.14\pm 0.38$\\
\\ \hline \\
\multirow{7}{4em}{500} &$HAL$\citep{chaudhry2020using}       &$41.79\pm 4.46$  &$84.54\pm 2.36$\\
             &$DER$ \citep{NEURIPS2020_b704ea2c}    &$70.51\pm 1.67$  &$93.40\pm 0.39$\\
             &$DER++$ \citep{NEURIPS2020_b704ea2c}  &$72.70\pm 1.17$  &$93.88\pm 0.60$\\
             &$Co^{2}L$ \citep{Cha_2021_ICCV}     &$\textbf{74.26}\pm \textbf{0.77}$  &$\textbf{95.90}\pm \textbf{0.26}$\\
             &$Ours$ $(\kappa =4)$                   &$71.51\pm 0.84$  &$95.37\pm 0.18$\\
             &$Ours$ $(\kappa =16)$                  &$71.68\pm 0.48$  &$95.51\pm 0.19$\\
             &$Ours$ $(\kappa =32)$                  &$71.32\pm 0.70$  &$95.43\pm 0.26$\\
\\ \hline \\
\end{tabular}
\end{center}
\end{table}
\subsection{Related works}
\paragraph{Contrastive Continual Learning}
Looking at continual learning from the perspective of feature learning, the former can be broken down into following two main questions: How can the learning features be robust and more reliable? How can the learned features be maintained? There are many recent works that focused on apply contrastive loss to leverage more generalized presentations. Self-supervised contrastive loss\citep{simclr} was adopted in these works \citep{Gallardo2021SelfSupervisedTE,DBLP:journals/corr/abs-2006-05882}, which increased models' performance without using labels. Supervised contrastive loss\citep{NEURIPS2020_d89a66c7} was used in many recent works\citep{Mai_2021_CVPR, Cha_2021_ICCV,han2021contrastive,Davari_2022_CVPR} to make samples that belong to the same class get closer to each other. Supervised contrastive loss was used in continual learning in the study presented by \citep{Mai_2021_CVPR}, then feature propagation was introduced with supervised contrastive loss to tackle multiple continual learning tasks\citep{han2021contrastive}. In \citep{Davari_2022_CVPR}, controlling feature forgetting was suggested using a linear classifier before and after a new task started. Another recent work\citep{Cha_2021_ICCV} improved supervised contrastive loss for better feature learning, and it also introduced a relation distillation to maintain the learned representation. There haven't been many works that consider a better approach on calculating similarity under the data imbalance environment. In our work, we investigate a different type of similarity function that adopts in asymmetric supervised contrastive loss\citep{Cha_2021_ICCV} to diminish intra-class variance.\\
\paragraph{von Mises-Fisher Distribution}
In recent years, von Mises–Fisher Distribution\citep{mardia_directional_1999} is one of the probability distributions on sphere that has been applied to many fields in deep learning. One of the previous works\citep{pmlr-v32-gopal14} presented a von Mises–Fisher distribution-based clustering method, which is more effective on high-dimensional data. Different types of von Mises–Fisher distribution based clustering approaches have been adopted for different tasks such as semantic segmentation\citep{hwang2019segsort} and graph neural network\citep{wang2023deep}. Using von Mises–Fisher distribution on metric learning or loss functions \citep{Hasnat2017vonMM,ZHE2019113,Kobayashi_2021_CVPR} showed significant improvement in performance as well.\\

\end{document}